# SeaFormer++: Squeeze-enhanced Axial Transformer for Mobile Visual Recognition

### v5

@@ 4.3 Ablation studies
Comparison with different self-attention modules
@@
Table 8: Results on ADE20K val set based on Swin Transformer architecture. (B) denotes backbone. OOM means CUDA out of memory. References: ISSA [66], A2-Nets [67]
To eliminate the impact of our architecture and demonstrate the effectiveness and generalization ability of SEA attention, we ran experiments on Swin Transformer [6] by replacing window attention in Swin Transformer with different attention blocks. We set the same training protocol, hyper-parameters, and model architecture configurations as Swin for a fair comparison. When replacing window attention with CCAttention (CCNet) or DoubleAttention (A2-Nets), they have much lower FLOPs than SeaFormer and other attention blocks. Considering that we may not be able to draw conclusions rigorously, we doubled the number of their Transformer blocks (including MLP). As ACmix has the same architecture configuration as Swin, we borrow the results from the original paper. From Table 8, it can be seen that SeaFormer outperforms other attention mechanisms with lower FLOPs and latency.
Effective and efficiency of SEA attention
@@
Table 7 summarizes the outcomes. Notably, the ‘Adaptive squeeze and Adaptive expand‘ method excelled, achieving 69.8% Top1 accuracy on ImageNet-1K and 36.8 mIoU on ADE20K. This method outperformed the ‘Mean pooling‘ and ‘Max pooling‘ methods combined with ‘Broadcast‘ expansion, both of which showed similar FLOPs and latency but lower performance metrics.
@@
We set four fusion methods, including ”Add directly”, ”Multiply directly”, ”Sigmoid add” and ”Sigmoid multiply”. X directly means features from context branch and spatial branch X directly. Sigmoid X means feature from context branch goes through a sigmoid layer and X feature from spatial branch.
From the Table 10 we can see that replacing sigmoid multiply with other fusion methods hurts performance. Sigmoid multiply is our optimal fusion block choice.
@@
M Params FLOPs Latency mIoU
@@
Posbias Params FLOPs Latency mIoU
@@
Table 11: Ablation studies on embedding dimensions and position bias.M = [128, 160] is an optimal embedding dimension in fusion blocks.
@@
Method Params FLOPs Top1 Latency
MobileNetV3-Small [31] 2.9M 0.1G 67.4 5ms
SeaFormer-T 1.8M 0.1G 67.9 7ms
SeaFormer-T++ 1.9M 0.1G 69.8 7ms
MobileViT-XXS [26] 1.3M 0.4G 69.0 24ms
MobileViTv2-0.50 [37] 1.4M 0.5G 70.2 32ms
MobileOne-S0* [68] 2.1M 0.3G 71.4 14ms
MobileNetV2 [30] 3.4M 0.3G 72.0 17ms
Mobile-Former96 [35] 4.8M 0.1G 72.8 31ms
SeaFormer-S 4.1M 0.2G 73.3 12ms
SeaFormer-S++ 4.2M 0.2G 74.5 12ms
EdgeViT-XXS [36] 4.1M 0.6G 74.4 71ms
LVT [34] 5.5M 0.9G 74.8 97ms
MobileViT-XS [26] 2.3M 0.9G 74.8 54ms
MobileNetV3-Large [31] 5.4M 0.2G 75.2 16ms
Mobile-Former151 [35] 7.7M 0.2G 75.2 42ms
MobileViTv2-0.75 [37] 2.9M 1.0G 75.6 68ms
MobileOne-S1* [68] 4.8M 0.8G 75.9 40ms
SeaFormer-B 8.7M 0.3G 76.0 20ms
SeaFormer-B++ 8.8M 0.3G 77.0 20ms
MobileOne-S2* [68] 7.8M 1.3G 77.4 63ms
EdgeViT-XS [36] 6.8M 1.1G 77.5 124ms
MobileViTv2-1.00 [37] 4.9M 1.8G 78.1 115ms
MobileOne-S3* [68] 10.1M 1.9G 78.1 91ms
MobileViT-S [26] 5.6M 1.8G 78.4 88ms
EfficientNet-B1 [54] 7.8M 0.7G 79.1 61ms
EfficientFormer-L1 [39] 12.3M 1.3G 79.2 94ms
Mobile-Former508 [35] 14.8M 0.5G 79.3 102ms
MobileOne-S4* [68] 14.8M 3.0G 79.4 143ms
SeaFormer-L 14.0M 1.2G 79.9 61ms
SeaFormer-L++ 14.1M 1.2G 80.6 61ms
Table 12: Image classification results on ImageNet-1K val set. The FLOPs and latency are measured with input size 224×224, except for MobileViT and MobileViTv2 that are measured with 256×256 according to their original implementations. * indicates re-parameterized variants [68]. The latency is measured on a single Qualcomm Snapdragon 865, and only an ARM CPU core is used for speed testing. No other means of acceleration, e.g., GPU or quantification, is used.
@@ 4.4 Image classification
We conduct experiments on ImageNet-1K [63], which contains 1.28M training images and 50K validation images from 1,000 classes. We employ an AdamW [69] optimizer for 600 epochs using a cosine decay learning rate scheduler. A batch size of 1024, an initial learning rate of 0.064, and a weight decay of 2e-5 are used. The results are illustrated in Table 12. Compared with other efficient approaches, SeaFormer++ achieves a relatively better trade-off between latency and accuracy.
(a) Squeeze Axial attention heatmaps
(b) Squeeze-enhanced Axial attention heatmaps
Figure 7: The visualization of attention heatmaps from the model consisting of squeeze Axial attention without detail enhancement (first row) and SeaFormer (second row). Heatmaps are produced by averaging channels of the features from the last attention block, normalizing to [0, 255] and up-sampling to the image size.
(a) Ground Truth
(b) TopFormer-B [32]
(c) SeaFormer-B (Ours)
Figure 8: Visualization of prediction results on ADE20K val set.
@@ 5 Object detection
We use RetinaNet [70] (one-stage) as the detection framework and follow the standard settings to use SeaFormer++ as backbone to generate e feature pyramid at multiple scales. All models are trained on train2017 split for 12 epochs (1×) from ImageNet pretrained weights.
From the table 13 We can observe that our SeaFormer++ achieves superior results on detection task which further demonstrates the strong generalization ability of our method.
Backbone AP FLOPs Params
ShuffleNetv2 [57] 25.9 161G 10.4M
SeaFormer-T 31.5 160G 10.9M
SeaFormer-T++ 32.8 160G 11.0M
Mobile-Former151 [35] 34.2 161G 14.4M
MobileNetV3 [31] 27.2 162G 12.3M
SeaFormer-S 34.6 161G 13.3M
SeaFormer-S++ 35.5 161G 13.4M
Mobile-Former214 [35] 35.8 162G 15.2M
Mobile-Former294 [35] 36.6 164G 16.1M
SeaFormer-B 36.7 164G 18.1M
SeaFormer-B++ 37.4 164G 18.2M
ResNet50 [56] 36.5 239G 37.7M
PVT-Tiny [71] 36.7 221G 23.0M
ConT-M [72] 37.9 217G 27.0M
SeaFormer-L 39.8 185G 24.0M
SeaFormer-L++ 40.2 185G 24.1M
Table 13: Results on COCO object detecion.
(a) Ground Truth
(b) SegFormer-B1 [4]
(c) SeaFormer-L (Ours)
Figure 9: Visualization of prediction results on ADE20K val set.
@@ 7.1 Attention heatmap
To demonstrate the effectiveness of detail enhancement in our squeeze-enhanced Axial attention (SEA attention), we ablate our model by removing the detail enhancement. We visualize the attention heatmaps of the two models in Figure 7. Without detail enhancement, attention heatmaps from solely SA attention appears to be axial strips while our proposed SEA attention is able to activate the semantic local region accurately, which is particularly significant in the dense prediction task.
@@ 7.2 Prediction results
We show the qualitative results and compare with the alternatives on the ADE20K validation set from two different perspectives. First we compare with a mobile-friendly rival TopFormer [32] with similar FLOPs and latency in Figure 8. Besides, we compare with the Transformer-based counterpart SegFormer-B1 [4] in Figure 9. In particular, our SeaFormer-L has lower computation cost than the SegFormer-B1. As shown in both figures, we demonstrate better segmentation results than both the mobile counterpart and Transformer-based approach.
@@ 8.1 Experimental setup
We perform multi-resolution distillation experiments over ADE20K [58]. The mean of intersection over union (mIoU) is set as the evaluation metric. We convert full-precision models to TNN [62] and measure latency on an ARM-based device with a single Qualcomm Snapdragon 865 processor. We set ADE20K fine-tuned network with original resolution input in section 4 as the teacher model and ImageNet-1K [63] pretrained network as the backbone of the student model. The input resolution of the student model is halved by average pooling.
The standard BatchNorm [64] layer is replaced by synchronized BatchNorm. The implementation of multi-resolution distillation is based on public codebase mmsegmentation [65]. We follow the batch size, training iteration scheduler and data augmentation strategy of TopFormer [32] and section 4 for a fair comparison. The initial learning rate is 0.0005 for batch size 32 or 0.00025 for batch size 16. The weight decay is 0.01. A “poly” learning rate scheduled with factor 1.0 is adopted. During inference, we set the same resize and crop rules as TopFormer to ensure fairness.
@@ 8.2 Comparison with state of the art
8.3 Ablation studie
This study conducts ablation experiments to evaluate various upsampling modules and loss function configurations for reducing computational demands and maintaining performance in visual tasks. It aims to identify efficient upsampling strategies and optimal loss combinations that preserve model accuracy. Experiments were standardized on the ADE20K validation set to ensure fair comparison and result reliability. The research compares the efficacy of techniques like bilinear interpolation, lightweight MobileNetV2-based upsampling, and standard convolutional upsampling, assessed by mean Intersection over Union (mIoU) and computational impact. Adjusting loss functions, including classification loss, cross-model classification loss, feature similarity loss, and output similarity loss, further analyzes their role in knowledge distillation effectiveness. These experiments offer insights into balancing efficiency and performance in model design, providing a methodology for exploring model enhancements under computational constraints and guiding optimal configurations for accurate visual recognition in practical applications.
Impact of upsample module design
@@ 8.3 Ablation studie
tandard convolution-based upsample module 2.7M 0.4G 35.7 30ms
@@
In this section, the effects of three upsampling strategies on model performance were compared. The direct interpolation approach, while requiring the least parameters (1.7M) and computational effort (0.3G FLOPs), resulted in the lowest mIoU (33.7%) and the least latency (20ms), suggesting limited complexity handling. The MobileNetV2-based lightweight upsampling improved mIoU to 35.5% with a slight latency increase to 22ms, offering a balanced performance enhancement. The standard convolution-based module, although yielding the highest mIoU (35.7%), did so at the cost of increased parameters (2.7M), computation (0.4G FLOPs), and latency (30ms). These findings highlight a trade-off between performance and speed in upsampling choices, with the MobileNetV2-based module providing an optimal balance for dense prediction tasks on resource-constrained devices.

### v6

@@ 4.3 Ablation studies
Comparing different self-attention modules in the Swin Transformer
@@
Table 8: Results on ADE20K val set based on Swin Transformer architecture. (B) denotes backbone. OOM means CUDA out of memory. References: ISSA Huang et al. (2019a), A2-Nets Chen et al. (2018b)
To eliminate the impact of our architecture and demonstrate the effectiveness and generalization ability of SEA attention, we ran experiments on Swin Transformer Liu et al. (2021) by replacing window attention in Swin Transformer with different attention blocks. We set the same training protocol, hyper-parameters, and model architecture configurations as Swin for a fair comparison. When replacing window attention with CCAttention (CCNet) or DoubleAttention (A2-Nets), they have much lower FLOPs than SeaFormer and other attention blocks. Considering that we may not be able to draw conclusions rigorously, we doubled the number of their Transformer blocks (including MLP). As ACmix has the same architecture configuration as Swin, we borrow the results from the original paper. From Table 8, it can be seen that SeaFormer outperforms other attention mechanisms with lower FLOPs and latency.
Comparing different self-attention modules in the SeaFormer
@@
Comparing different architecture with the same attention module
Table 10 provides a comprehensive comparison of different architectures with the same attention module. Our proposed architecture, SeaFormer++, demonstrates a favorable balance between efficiency and accuracy. Notably, SeaFormer++ achieves a latency of 369ms with only 6.5G FLOPs, which is significantly lower than Swin Transformer (2278ms and 24.9G FLOPs) while maintaining competitive mIoU performance (43.8 vs. 46.5).
Additionally, with the inclusion of knowledge distillation (SeaFormer++ KD), the model’s latency further improves to an impressive 177ms with only 2.9G FLOPs, while still maintaining a mIoU of 42.2. This demonstrates the impact of knowledge distillation in not only reducing computational cost but also preserving performance. Compared to DeepLabV3+ and SegFormer, our SeaFormer++ with KD achieves comparable accuracy (42.2 mIoU vs. 42.7 mIoU for SegFormer) while offering much faster inference times and fewer FLOPs, confirming the efficiency of the overall architecture design and distillation pipeline.
Architecture Params FLOPs Latency mIoU
DeepLabV3+ Chen et al. (2018a) 17.1M 24.5G 710ms 39.3
SegFormer Xie et al. (2021) 17.1M 15.2G 920ms 42.7
Swin Liu et al. (2021) 34.0M 24.9G 2278ms 46.5
SeaFormer++ 14.1M 6.5G 369ms 43.8
SeaFormer++(KD) 17.1M 2.9G 177ms 42.2
Table 10: Comparison of different architecture with our proposed SEA attention.
@@
Table 7 summarizes the outcomes. Notably, the ‘Adaptive squeeze and Adaptive expand’ method excelled, achieving 69.8% Top1 accuracy on ImageNet-1K and 36.8 mIoU on ADE20K. The 1x1 convolutional layers in the adaptive squeeze and expand modules contribute a minor increase in parameter count (+0.1M) and computational cost (+0.01G), which have a negligible impact on latency (+1ms). However, these modifications result in significant performance improvement, demonstrating the effectiveness of our design choices.
@@
We set four fusion methods, including Add directly, Multiply directly, Sigmoid add and Sigmoid multiply. X directly means features from context branch and spatial branch X directly. Sigmoid X means feature from context branch goes through a sigmoid layer and X feature from spatial branch.
From Table 11 we can see that replacing sigmoid multiply with other fusion methods hurts performance. Sigmoid multiply is our optimal fusion block choice.
@@
Embed dim Params FLOPs Latency mIoU
@@
Position bias Params FLOPs Latency mIoU
@@
Table 12: Ablation studies on embedding dimensions and position bias. [128, 160] is an optimal embedding dimension in fusion blocks.
@@
Model Head Params FLOPs Latency mIoU
Topformer-Tiny* Simple head 1.4M 0.58G 43ms 34.6
Topformer-Tiny* Light head 1.4M 0.64G 47ms 33.1
SeaFormer-Tiny* Simple head 1.7M 0.60G 42ms 35.6
SeaFormer-Tiny* Light head 1.7M 0.60G 40ms 35.8
SeaFormer-Tiny++ Light head 1.8M 0.61G 41ms 36.8
Topformer-Small* Simple head 3.1M 1.16G 74ms 37.0
Topformer-Small* Light head 3.1M 1.19G 76ms 35.3
SeaFormer-Small* Simple head 4.0M 1.10G 70ms 38.8
SeaFormer-Small* Light head 4.0M 1.10G 67ms 39.4
SeaFormer-Small++ Light head 4.1M 1.11G 68ms 39.7
Topformer-Base* Simple head 5.1M 1.81G 110ms 39.2
Topformer-Base* Light head 5.1M 1.85G 112ms 36.6
SeaFormer-Base* Simple head 8.7M 1.84G 108ms 40.6
SeaFormer-Base* Light head 8.6M 1.82G 106ms 41.0
SeaFormer-Base++ Light head 8.7M 1.83G 107ms 41.4
Table 13: Comparison of different heads and backbones across various models. * indicates training batch size is 32.
Additional ablation studies
To justify the effectiveness of the backbone and decoder in SeaFormer++, we have conducted additional experiments comparing different combinations of backbones and heads. The results are shown in Table 13.
We used the same backbone with different heads, including the Simple head from the most related baseline Topformer Zhang et al. (2022b) and the Light head proposed in our work. Compared to the Simple head in SeaFormer-Base, our proposed Light head gains 0.4 mIoU improvement with a decrease of 0.1M parameters, 0.02G FLOPs, and 2ms Latency. Additionally, we compared different backbones (Topformer, SeaFormer, and SeaFormer++) using the same proposed Light head, SeaFormer-Base++ outperforms TopFormer-Base (41.4 vs.36.6) with lower latency (107ms vs.112ms).
The results in the table highlight the superior performance of the SeaFormer++ backbone and Light head in terms of both efficiency (latency and FLOPs) and effectiveness (mIoU). SeaFormer++ consistently outperforms Topformer across various configurations, confirming the strength of our backbone and head design.
@@ 4.4 Image classification
We conduct experiments on ImageNet-1K Deng et al. (2009), which contains 1.28M training images and 50K validation images from 1,000 classes. We employ an AdamW Kingma (2014) optimizer for 600 epochs using a cosine decay learning rate scheduler. A batch size of 1024, an initial learning rate of 0.064, and a weight decay of 2e-5 are used. The results are illustrated in Table 14. Compared with other efficient approaches, SeaFormer++ achieves a relatively better trade-off between latency and accuracy.
@@ 5 Object detection
We use RetinaNet Lin (2017) (one-stage) as the detection framework and follow the standard settings to use SeaFormer++ as backbone to generate e feature pyramid at multiple scales. All models are trained on train2017 split for 12 epochs (1×) from ImageNet pretrained weights.
From the table 15, we can observe that our SeaFormer++ achieves superior results on detection task which further demonstrates the strong generalization ability of our method.
@@ 6 Performance under different precision of the models
Method Params FLOPs Top1 Latency
MobileNetV3-Small Howard et al. (2019) 2.9M 0.1G 67.4 5ms
SeaFormer-T 1.8M 0.1G 67.9 7ms
SeaFormer-T++ 1.9M 0.1G 69.8 7ms
MobileViT-XXS Mehta and Rastegari (2022a) 1.3M 0.4G 69.0 24ms
MobileViTv2-0.50 Mehta and Rastegari (2022b) 1.4M 0.5G 70.2 32ms
MobileOne-S0* Vasu et al. (2023) 2.1M 0.3G 71.4 14ms
MobileNetV2 Sandler et al. (2018) 3.4M 0.3G 72.0 17ms
Mobile-Former96 Chen et al. (2022b) 4.8M 0.1G 72.8 31ms
SeaFormer-S 4.1M 0.2G 73.3 12ms
SeaFormer-S++ 4.2M 0.2G 74.5 12ms
EdgeViT-XXS Pan et al. (2022a) 4.1M 0.6G 74.4 71ms
LVT Yang et al. (2022a) 5.5M 0.9G 74.8 97ms
MobileViT-XS Mehta and Rastegari (2022a) 2.3M 0.9G 74.8 54ms
MobileNetV3-Large Howard et al. (2019) 5.4M 0.2G 75.2 16ms
Mobile-Former151 Chen et al. (2022b) 7.7M 0.2G 75.2 42ms
MobileViTv2-0.75 Mehta and Rastegari (2022b) 2.9M 1.0G 75.6 68ms
EfficientFormerV2-S0 Li et al. (2023b) 3.5M 0.4G 75.7 25ms
MobileOne-S1* Vasu et al. (2023) 4.8M 0.8G 75.9 40ms
SeaFormer-B 8.7M 0.3G 76.0 20ms
SeaFormer-B++ 8.8M 0.3G 77.0 20ms
MobileOne-S2* Vasu et al. (2023) 7.8M 1.3G 77.4 63ms
EdgeViT-XS Pan et al. (2022a) 6.8M 1.1G 77.5 124ms
MobileViTv2-1.00 Mehta and Rastegari (2022b) 4.9M 1.8G 78.1 115ms
MobileOne-S3* Vasu et al. (2023) 10.1M 1.9G 78.1 91ms
AxialFormer Wang et al. (2020a) 11.6M 3.0G 78.1 151ms
MobileViT-S Mehta and Rastegari (2022a) 5.6M 1.8G 78.4 88ms
EfficientNet-B1 Tan and Le (2019) 7.8M 0.7G 79.1 61ms
EfficientFormer-L1 Li et al. (2022) 12.3M 1.3G 79.2 94ms
Mobile-Former508 Chen et al. (2022b) 14.8M 0.5G 79.3 102ms
MobileOne-S4* Vasu et al. (2023) 14.8M 3.0G 79.4 143ms
SeaFormer-L 14.0M 1.2G 79.9 61ms
SeaFormer-L++ 14.1M 1.2G 80.6 61ms
Table 14: Image classification results on ImageNet-1K val set. The FLOPs and latency are measured with input size 224×224, except for MobileViT and MobileViTv2 which are measured with 256×256 according to their original implementations. * indicates re-parameterized variants Vasu et al. (2023). The latency is measured on a single Qualcomm Snapdragon 865, and only an ARM CPU core is used for speed testing. No other means of acceleration, e.g., GPU or quantification, is used.
(a) Squeeze Axial attention heatmaps
(b) Squeeze-enhanced Axial attention heatmaps
Figure 7: The visualization of attention heatmaps from the model consisting of squeeze axial attention without detail enhancement (first row) and SeaFormer (second row). Heatmaps are produced by averaging channels of the features from the last attention block, normalizing to [0, 255], and up-sampling to the image size.
(a) Ground Truth
(b) TopFormer-B Zhang et al. (2022b)
(c) SeaFormer-B (Ours)
Figure 8: Visualization of prediction results on ADE20K val set.
Backbone AP FLOPs Params
ShuffleNetv2 25.9 161G 10.4M
SeaFormer-T 31.5 160G 10.9M
SeaFormer-T++ 32.8 160G 11.0M
Mobile-Former151 34.2 161G 14.4M
MobileNetV3 27.2 162G 12.3M
SeaFormer-S 34.6 161G 13.3M
SeaFormer-S++ 35.5 161G 13.4M
Mobile-Former214 35.8 162G 15.2M
Mobile-Former294 36.6 164G 16.1M
SeaFormer-B 36.7 164G 18.1M
SeaFormer-B++ 37.4 164G 18.2M
ResNet50 36.5 239G 37.7M
PVT-Tiny 36.7 221G 23.0M
AxialFormer 37.7 210G 23.7M
ConT-M 37.9 217G 27.0M
SeaFormer-L 39.8 185G 24.0M
SeaFormer-L++ 40.2 185G 24.1M
Table 15: Results on COCO object detection. References: ShuffleNetv2 Ma et al. (2018), Mobile-Former151 Chen et al. (2022b), MobileNetV3 Howard et al. (2019), ResNet50 He et al. (2016), PVT-Tiny Wang et al. (2021), AxialFormer Wang et al. (2020a), ConT-M Yan et al. (2021).
(a) Ground Truth
(b) SegFormer-B1 Xie et al. (2021)
(c) SeaFormer-L (Ours)
Figure 9: Visualization of prediction results on ADE20K val set.
Figure 10: The inference latency of components.
@@ 7.1 Attention heatmap
To demonstrate the effectiveness of detail enhancement in our squeeze-enhanced Axial attention (SEA attention), we ablate our model by removing the detail enhancement. We visualize the attention heatmaps of the two models in Figure 7. Without detail enhancement, attention heatmaps from solely SA attention appear to be axial strips while our proposed SEA attention is able to activate the semantic local region accurately, which is particularly significant in the dense prediction task.
@@ 7.2 Prediction results
We show the qualitative results and compare with the alternatives on the ADE20K validation set from two different perspectives. First, we compare with a mobile-friendly rival TopFormer Zhang et al. (2022b) with similar FLOPs and latency in Figure 8. Besides, we compare with the Transformer-based counterpart SegFormer-B1 Xie et al. (2021) in Figure 9. In particular, our SeaFormer-L has a lower computation cost than the SegFormer-B1. As shown in both figures, we demonstrate better segmentation results than both the mobile counterpart and Transformer-based approach.
@@ 8.1 Experimental setup
We perform multi-resolution distillation experiments over ADE20K Zhou et al. (2017). mIoU is set as the evaluation metric. We convert full-precision models to TNN Contributors (2019) and measure latency on an ARM-based device with a single Qualcomm Snapdragon 865 processor. We set ADE20K fine-tuned network with original resolution input in section 4 as the teacher model and ImageNet-1K Deng et al. (2009) pretrained network as the backbone of the student model. The input resolution of the student model is halved by average pooling.
The standard BatchNorm Ioffe and Szegedy (2015) layer is replaced by synchronized BatchNorm. The implementation of multi-resolution distillation is based on public codebase mmsegmentation Contributors (2020). We follow the batch size, training iteration scheduler and data augmentation strategy of TopFormer Zhang et al. (2022b) and section 4 for a fair comparison. The initial learning rate is 0.0005 for batch size 32 or 0.00025 for batch size 16. The weight decay is 0.01. A poly learning rate scheduled with factor 1.0 is adopted. During inference, we set the same resize and crop rules as TopFormer to ensure fairness.
@@ 8.2 Comparison with state of the art
Standard convolution-based upsample module 2.7M 0.4G 35.7 30ms
@@
Teacher resolution Cls loss Out loss Feat loss Cross loss mIoU
- 32.5
512x512 33.7
512x512 34.7
512x512 35.5
256x256 32.1
Table 18: Ablation study results of loss function design. The resolution of the student model is 256x256.
8.3 Ablation studies
This study conducts ablation experiments to evaluate various upsampling modules and loss function configurations for reducing computational demands and maintaining performance in visual tasks. It aims to identify efficient upsampling strategies and optimal loss combinations that preserve model accuracy. Experiments were standardized on the ADE20K validation set to ensure fair comparison and result reliability. The research compares the efficacy of techniques like bilinear interpolation, lightweight MobileNetV2-based upsampling, and standard convolutional upsampling, assessed by mIoU and computational impact. Adjusting loss functions, including classification loss, cross-model classification loss, feature similarity loss, and output similarity loss, further analyzes their role in knowledge distillation effectiveness. These experiments offer insights into balancing efficiency and performance in model design, providing a methodology for exploring model enhancements under computational constraints and guiding optimal configurations for accurate visual recognition in practical applications.
Impact of upsample module design
In this section, the effects of three upsampling strategies on model performance were compared. The direct interpolation approach, while requiring the least parameters (1.7M) and computational effort (0.3G FLOPs), resulted in the lowest mIoU (33.7%) and the least latency (20ms), suggesting limited complexity handling. The MobileNetV2-based lightweight upsampling improved mIoU to 35.5% with a slight latency increase to 22ms, offering a balanced performance enhancement. The main components in the MobileNetV2-based upsampling module are 1x1 convolutional layers. From the table we conclude that 1x1 convolutional layers in the upsampling module contribute a minor increase in parameter count (+0.6M), which has a negligible impact on latency (+2ms) with a significant performance boost, from mIoU 33.7 to mIoU 35.5, demonstrating the effectiveness of our model design. The standard convolution-based module, although yielding the highest mIoU (35.7%), did so at the cost of increased parameters (2.7M), computation (0.4G FLOPs), and latency (30ms). These findings highlight a trade-off between performance and speed in upsampling choices, with the MobileNetV2-based module providing an optimal balance for dense prediction tasks on resource-constrained devices.